# Permutation twisted cohomology, remixed

### v2

@@ A $kG$-module $M$ is:
    a $p$-permutation module (where $p$ denotes the characteristic of $k$) if $M$ is a direct summand of a permutation module, or equivalently, $\operatorname{Res}^{G}_{S}M$ is a permutation module for some Sylow $p$-subgroup $S$ of $G$. In particular, if $G$ is a finite $p$-group, every $p$-permutation module is a permutation module.
We set $\mathscr{K}(G):=\operatorname{K}_{b}(p\operatorname{-perm})$, the bounded homotopy category of $p$-permutation modules.
@@
Let $K\leq H\leq G$ be subgroups of $G$. The augmentation homomorphism for $G$-sets $\operatorname{aug}\colon G/K\to G/H$ is the homomorphism induced by the assignment $1K\mapsto 1H$. The augmentation homomorphism between the permutation modules $\operatorname{aug}\colon k[G/K]\to k[G/H]$ is the induced map on vector spaces. We note the standard example of the augmentation homomorphism is when $H=G$ and $K=1$.
@@
Remark 1.3.
By [BG23a], we have a surjective tt-functor $\mathscr{K}(G)\twoheadrightarrow\operatorname{D}_{b}(kG)$ with kernel $\mathscr{K}_{ac}(G)$, the tt-ideal consisting of all acyclic complexes. Set $V_{G}:=\operatorname{Spc}(\operatorname{D}_{b}(kG))$, then [BG25, Proposition 3.22] asserts that this functor induces an open inclusion $V_{G}\hookrightarrow\operatorname{Spc}(\mathscr{K}(G))$. We say $V_{G}\subseteq\operatorname{Spc}(\mathscr{K}(G))$ is the cohomological open. As noted in the introduction of [BG25], the crux of the matter is its closed complement, $\operatorname{supp}(\mathscr{K}_{ac})$.
The Brauer quotient is an explicit construction on the level of modules which mimics taking fixed points of $G$-sets. Balmer-Gallauer adapt this construction to work on the level of big categories, dubbing them modular fixed points functors.
@@
Notation 1.5.
@@
In fact, the Balmer spectrum, as a set, is built entirely from these “$p$-local” cohomological opens.
@@
[BG25, Theorem 2.10] Every point of $\operatorname{Spc}(\mathscr{K}(G))$ is the image $\hat{\psi}^{H}(\mathfrak{p})$ of a point $\mathfrak{p}\in V_{{G}/\!\!/{H}}$ for some subgroup $H\leq G$ unique up to $G$-conjugation, i.e. we have $\hat{\psi}^{H}(\mathfrak{p})=\hat{\psi}^{H^{\prime}}(\mathfrak{p}^{\prime})$ if and only if there exists a $g\in G$ such that $^{g}H=H^{\prime}$ and $^{g}\mathfrak{p}=\mathfrak{p}^{\prime}$.
Notation 1.7.
We denote by $\mathscr{P}_{G}(H,\mathfrak{p})\in\operatorname{Spc}(\mathscr{K})$ the image of the prime $\mathfrak{p}\in V_{{G}/\!\!/{H}}$ under $\hat{\psi}^{H}:V_{{G}/\!\!/{H}}\to\operatorname{Spc}(\mathscr{K}(G))$. By [BG25, Theorem 2.10], every prime is, up to $G$-conjugation, uniquely expressible in this way.
Theorem 1.8.
[BG25, Theorem 2.11] The family of functors
indexed by conjugacy classes of $p$-subgroups $H\leq G$ is conservative. This restricts to a conservative family of functors $\{\mathscr{K}(G)\xrightarrow{\Psi^{H}}\mathscr{K}({G}/\!\!/{H})\twoheadrightarrow\operatorname{D}_{b}(k[{G}/\!\!/{H}])\}_{H\in s_{p}(G)/G}$ on compacts.
Next, we briefly review the classification of endotrivial complexes.
@@
A chain complex $C\in\mathscr{K}(G)$ is endotrivial if it is an invertible object, i.e. $C^{*}\otimes_{k}C\simeq k[0]$.
Remark 1.10.
By the Künneth formula for $kG$-modules, for any $p$-subgroup $H\leq G$, $\Psi^{H}(C)$ has nonzero homology in exactly one homological degree, with that homology having $k$-dimension one. Let $h_{C}(H)\in\mathbb{Z}$ denote the unique integer for which $\operatorname{H}_{h_{C}(H)}(\Psi^{H}(C))\neq 0$.
@@
    The function $h_{C}\colon s_{p}(G)\to\mathbb{Z}$ is a well-defined superclass function, i.e. a function constant on conjugacy classes of subgroups of $G$.
@@
    The assignment $h\colon\operatorname{Pic}(\mathscr{K}(G))\to\operatorname{CF}(G,p)$ given by $C\mapsto h_{C}$ is a well-defined group homomorphism.
@@
The collection of Borel-Smith functions forms an additive subgroup $\operatorname{CF}_{b}(G,p)$ of $\operatorname{CF}(G,p)$. In fact, if $G$ is a finite $p$-group, then under the identification $\operatorname{CF}(G)\cong B^{*}(G):=\operatorname{Hom}(B(G),\mathbb{Z})$, $\operatorname{CF}_{b}(G)$ forms a rational $p$-biset subfunctor of $\operatorname{CF}(G)$. See [BY07, Proposition 3.7] for details.
@@
    [Mil25a, Theorem 4.6] Let $G$ be a finite $p$-group. The h-mark homomorphism is an isomorphism onto the subgroup of Borel-Smith functions
@@
    [Mil25a, Corollary 6.4] Let $G$ be a finite group. The h-mark homomorphism has image $\operatorname{CF}_{b}(G,p)$ and induces a split exact sequence
@@
Let $G$ be a finite $p$-group. The first nontrivial example of an endotrivial, and those which make the magic happen in the elementary abelian case, are as follows (c.f. [BG25, Definition 12.3]). If $p=2$, one has an endotrivial for $C_{2}$ given by $kC_{2}\to k,$ with $k$ in homological degree 0 and the nonzero differential given by the augmentation homomorphism. Otherwise if $p$ is odd, one has an endotrivial for $C_{p}$ by truncating a periodic resolution of the trivial $kC_{p}$-module $k$, $kC_{p}\to kC_{p}\to k$ with $k$ in homological degree 0. For any subgroup $N\triangleleft G$ of index $p$, inflation yields the following complex
@@
Here $p$ is assumed odd, if $p=2$ then the highest degree term can be deleted.
A less easy example of an endotrivial for $p=2$ and $G=D_{2^{n}}$ is as follows. Let $H_{1},H_{2}$ be nonconjugate, noncentral subgroups of order 2 (this choice is unique up to conjugacy and reordering). Then the following complex is endotrivial, and a small example of a nontrivial faithful, irreducible endotrivial, i.e. one arising from a faithful irreducible real representation (see Definition 2.5),
$kD_{2^{n}}\to k[D_{2^{n}}/H_{1}]\oplus k[D_{2^{n}}/H_{1}]\to k,$
where all maps between indecomposable permutation modules are augmentation homomorphisms.
2. Borel-Smith functions and representation spheres
We begin by considering the topological properties of the main characters of our story, the endotrivials arising from genuine real representations, i.e. representations spheres. Such endotrivials come from objects of higher structure; they arise from chain complexes of free modules over the orbit category $\Gamma_{G}$, and in particular, their differentials are augmentation homomorphisms which are well-behaved with respect to local-global considerations. We discuss this now.
@@ 2. Borel-Smith functions and representation spheres
This choice of terminology will become clear in the sequel. We first recall some important facts about Borel-Smith functions. Let $G$ be a nilpotent group and $F$ a field. Given a $FG$-module $V$, the dimension function associated to $V$ is the superclass function
@@
If $F$ has characteristic 0, this induces a group homomorphism $R_{F}(G)\to\operatorname{CF}(G)$. When $F=\mathbb{R}$, topologists may write $RO(G)=R_{\mathbb{R}}(G)$, as a representation sphere is nothing more than a one-point compactification of a real representation.
@@
[tD87, Theorem 5.4 and Theorem 5.13, pages 211 and 216] Let $G$ be a nilpotent group. The image of $\dim\colon RO(G)\to\operatorname{CF}(G)$ is the group of Borel-Smith functions $\operatorname{CF}_{b}(G)$. Moreover, if $f$ is a effective Borel-Smith function, there exists a real representation $V$ for which $\dim(V)=f$.
@@
Proof.
This follows from [tD87, Theorem 5.4, page 211] since the dimension function of any representation is effective, and [Mil25a, Theorem 4.6], since the image of the h-mark homomorphism is precisely the group of Borel-Smith functions of $G$. ∎
The previous corollary tells us that it suffices to work with effective endotrivials, which are particularly well-behaved. In fact, when $G$ is a nilpotent group, the dimension function gives a canonical effective basis of $\operatorname{CF}_{b}(G)$, which arises from the real irreps of $G$. Therefore, if $G$ is a finite $p$-group, we have a corresponding canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K})$ consisting of effective endotrivials.
@@
Since the image of the dimension homomorphism is precisely $\operatorname{CF}_{b}(G)$, it suffices to show that the set $\mathcal{B}:=\{f_{1},\dots,f_{n}\}$ is linearly independent after removing duplicates. [tD87, Proposition 5.9, page 213] asserts that $\ker(\dim)$ is generated by elements of the form $V-\psi^{k}(V)$, where $V$ is an irreducible real representation, $\psi^{k}$ is the $k$-th Adams operation, and $k$ is coprime to $|G|$. Therefore, the duplicates in $\mathcal{B}$ arise from Adams operation conjugates, and it follows that any set of real irreducible representations which are not Adams operation conjugates will correspond to a linearly independent set of Borel-Smith functions. ∎
@@
We call the Borel-Smith functions (resp. endotrivials) associated to the real irreps of $G$ the irreducible Borel-Smith functions (resp. endotrivials). These objects are necessarily effective.
@@
$\dim_{\mathbb{R}}V^{H}=\frac{1}{|H|}\sum_{h\in H}\chi(h),$
a practical method of computing the basis of $\operatorname{CF}_{b}(G)$.
Checking character tables of $G$ a finite $p$-group of normal $p$-rank one shows the basis of $\operatorname{Pic}(\mathscr{K}(G))$ obtained in [Mil24, Section 6] coincides with the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$. (The author remarks that he finds this is rather surprising, as the computations of [Mil24] were performed entirely ad-hoc and were done prior to the classification of $\operatorname{Pic}(\mathscr{K}(G))$.)
Notation 2.7.
Let $G$ be a finite $p$-group. Given a $\mathbb{R}G$-module $V$, let $f_{V}$ and $C_{V}$ denote the corresponding effective Borel-Smith function and effective indecomposable endotrivial respectively. That is, $f_{V}:=\dim V$ and $C_{V}$ is the unique indecomposable endotrivial for which $h_{C_{V}}=\dim V$.
Per tom Dieck [tD87], Borel-Smith functions correspond to representation spheres, which is simply a manifestation of 2.2. If $X$ is a finite-dimensional $G$-complex which is a $\mathbb{F}_{p}$-cohomology sphere, the fixed-point set $X^{H}$ is a $\mathbb{F}_{p}$-cohomology-sphere of dimension $d(H)$. The assignment $H\mapsto d(H)+1$ is (unfortunately) also called the dimension function of $X$. The set of dimension functions of representation spheres is precisely the effective non-negative Borel-Smith functions. Given a real representation $V$, let $S^{V}$ denote the corresponding representation sphere.

### v3

@@ A $kG$-module $M$ is:
    a $p$-permutation module (where $p$ denotes the characteristic of $k$) if $M$ is a direct summand of a permutation module, or equivalently, $\operatorname{Res}^{G}_{S}M$ is a permutation module for some Sylow $p$-subgroup $S$ of $G$.
    In particular, if $G$ is a finite $p$-group, every $p$-permutation module is a permutation module. In this case, the indecomposable permutation modules are precisely those isomorphic to $kX$ for $X$ a transitive $G$-set. If $G$ is not a $p$-group, these modules may decompose further.
We set $\mathscr{K}(G):=\operatorname{K}_{b}(p\operatorname{-perm}(kG))$, the bounded homotopy category of $p$-permutation modules. The category $\mathscr{K}(G)$ is a rigid tensor-triangulated category, i.e. there is a given by the internal hom $C^{*}:=\operatorname{Hom}_{k}(C,k)$. It is an essentially small and is the compact part of a rigidly-compactly-generated (i.e., “big”) tensor-triangulated category $\mathscr{D}(G)$, the derived category of permutation $kG$-modules. It
We note this category can be defined in the context of cohomological Mackey functors, Artin motives, or equivariant stable homotopy theory, and refer the reader to [BG23b, FUH25] for details on these equivalences.
There are two standard homomorphisms between permutation modules we’ll frequently consider.
@@
Let $K\leq H\leq G$ be subgroups of $G$. The augmentation homomorphism between the permutation modules $\operatorname{aug}\colon k[G/K]\to k[G/H]$ is the homomorphism induced by the $G$-set morphism given by $1K\mapsto 1H$. The standard example of the augmentation homomorphism is when $H=G$ and $K=1$.
@@
Definition 1.6.
By [BG23a], we have a localization $\mathscr{K}(G)\twoheadrightarrow\mathscr{K}(G)/\mathscr{K}_{ac}(G)\cong\operatorname{D}_{b}(kG)$, where $\mathscr{K}_{ac}(G)$ denotes the thick $\otimes$-ideal of $\mathscr{K}(G)$ consisting of all acyclic complexes. We set $V_{G}:=\operatorname{Spc}(\operatorname{D}_{b}(kG))$, then [BG25, Proposition 3.22] asserts that this functor induces an open inclusion $V_{G}\hookrightarrow\operatorname{Spc}(\mathscr{K}(G))$. The subset $V_{G}\subseteq\operatorname{Spc}(\mathscr{K}(G))$ is called the cohomological open. As noted in the introduction of [BG25], the “interesting” part of $\mathscr{K}(G)$ is its closed complement, $\operatorname{supp}(\mathscr{K}_{ac})$.
Brauer quotients are explicit constructions on $kG$-modules which mimic taking $H$-fixed points of $G$-sets when $H$ is a $p$-subgroup of $G$. Balmer–Gallauer adapt this construction to work on the level of big categories, dubbing them modular fixed points functors.
@@
This conservativity theorem is an example of “local-to-global” phenomena in modular representation theory.
Theorem 1.8.
[BG25, Theorem 2.11] The family of functors
indexed by conjugacy classes of $p$-subgroups $H\leq G$ is conservative, i.e., detects vanishing of objects. This restricts to a conservative family of functors
$\{\mathscr{K}(G)\xrightarrow{\Psi^{H}}\mathscr{K}({G}/\!\!/{H})\twoheadrightarrow\operatorname{D}_{b}(k[{G}/\!\!/{H}])\}_{H\in s_{p}(G)/G}$
on the compact parts.
Notation 1.9.
@@
These modular fixed point functors and the localization to the bounded derived category are the keys to deducing $\operatorname{Spc}(\mathscr{K}(G))$ as a set. The Balmer spectrum, as a set, is built entirely from these $p$-local cohomological opens.
@@
[BG25, Theorem 2.10] Every point of $\operatorname{Spc}(\mathscr{K}(G))$ is the image $\hat{\psi}^{H}(\mathfrak{p})$ of a point $\mathfrak{p}\in V_{{G}/\!\!/{H}}$ for some subgroup $H\leq G$ unique up to $G$-conjugation, i.e., we have $\hat{\psi}^{H}(\mathfrak{p})=\hat{\psi}^{H^{\prime}}(\mathfrak{p}^{\prime})$ if and only if there exists a $g\in G$ such that $^{g}H=H^{\prime}$ and $^{g}\mathfrak{p}=\mathfrak{p}^{\prime}$. In other words,
$\operatorname{Spc}(\mathscr{K}(G))=\bigsqcup_{H\in s_{p}(G)/G}\operatorname{Spc}(\operatorname{D}_{b}(k[{G}/\!\!/{H}]))\cong\bigsqcup_{H\in s_{p}(G)/G}\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H};k)).$
Notation 1.11.
We denote by $\mathscr{P}_{G}(H,\mathfrak{p})\in\operatorname{Spc}(\mathscr{K}(G))$ the image of the prime $\mathfrak{p}\in V_{{G}/\!\!/{H}}$ under $\hat{\psi}^{H}:V_{{G}/\!\!/{H}}\to\operatorname{Spc}(\mathscr{K}(G))$. We omit the $G$ subscript if the ambient group is clear. By [BG25, Theorem 2.10], every prime is, up to $G$-conjugation, uniquely expressible in this way.
The subset $\{\mathscr{P}(H,\mathfrak{p})\mid\mathfrak{p}\in V_{{G}/\!\!/{H}}\}\subseteq\operatorname{Spc}(\mathscr{K}(G))$ is generally not open nor closed when $H\neq 1$, but contains the closed point $\mathscr{P}(H,0)\in\operatorname{Spc}(\mathscr{K}(G))$. Explicitly, one can have an inclusion of primes $\mathscr{P}(H,\mathfrak{p})\subseteq\mathscr{P}(H^{\prime},\mathfrak{q})$ where $H>_{G}H^{\prime}$, see [BG25, Corollary 7.13].
2. Preliminaries: Endotrivial complexes and Borel-Smith functions
We next review the classification of endotrivial complexes, following [MIL25b].
@@ 2. Preliminaries: Endotrivial complexes and Borel-Smith functions
A chain complex $C\in\mathscr{K}(G)$ is endotrivial if it is an invertible object in $\mathscr{K}(G)$, i.e., we have $C^{*}\otimes C\cong k$. The set $\operatorname{Pic}(\mathscr{K}(G))$ of all endotrivials up to isomorphism forms a group under $\otimes$, and is called the Picard group of $\mathscr{K}(G)$.
By the Künneth formula, for any $p$-subgroup $H\leq G$, $\Psi^{H}(C)$ has nonzero homology in exactly one homological degree, with that homology having $k$-dimension one. Let $h_{C}(H)\in\mathbb{Z}$ denote the unique integer for which $\operatorname{H}_{h_{C}(H)}(\Psi^{H}(C))\neq 0$. The next proposition essentially reduces the study of endotrivials to considering superclass functions.
@@
    The function $h_{C}\colon s_{p}(G)\to\mathbb{Z}$ is a well-defined superclass function.
@@
    The assignment $h\colon\operatorname{Pic}(\mathscr{K}(G))\to\operatorname{CF}(G,p)$ given by $C\mapsto h_{C}$ is a well-defined group homomorphism, with kernel isomorphic to $\operatorname{Hom}(G,k^{\times})$.
Definition 2.3.
@@
By [MIL24, Theorem 3.6], to determine $\operatorname{Pic}(\mathscr{K}(G))$, it effectively suffices to compute $\operatorname{im}(h)$.
@@
The collection of Borel-Smith functions forms an additive subgroup $\operatorname{CF}_{b}(G,p)$ of $\operatorname{CF}(G,p)$. In fact, if $G$ is a finite $p$-group, then under the identification $\operatorname{CF}(G)\cong B^{*}(G):=\operatorname{Hom}(B(G),\mathbb{Z})$, where $B(G)$ denotes the Burnside ring of $G$, $\operatorname{CF}_{b}(G)$ is a rational $p$-biset subfunctor of $\operatorname{CF}(G)$. See [BY07, Proposition 3.7] for details.
@@
    [MIL25b, Theorem 4.6] Let $G$ be a finite $p$-group. The h-mark homomorphism is an isomorphism onto the subgroup of Borel-Smith functions
@@
    [MIL25b, Corollary 6.4] Let $G$ be a finite group. The h-mark homomorphism has image $\operatorname{CF}_{b}(G,p)$ and induces a split exact sequence
@@
Remark 2.6.
The classification of endotrivials for a $p$-group can be deduced in multiple ways; one way is by identifying them with representation spheres, which we discuss in the following section. This method was considered by Bachmann in his dissertation [BAC16]. Another way, the method we used in [MIL25b], is by recovering a short exact sequence constructed by Bouc–Yalçin [BY07] relating Borel-Smith functions to the Dade group of a finite group. This group parametrizes endopermutation $kG$-modules, modules $M$ for which $M^{*}\otimes M$ is permutation. The classification of these modules for $p$-groups was completed by Bouc [BOU06].
The classification of endotrivials for a non-$p$-group follows from some technical results concerning the homomorphism induced by restriction $\operatorname{Res}^{G}_{H}\colon\operatorname{Pic}(\mathscr{K}(G))\to\operatorname{Pic}(\mathscr{K}(H))$ for subgroups $H$ containing a Sylow $p$-subgroup of $G$ deduced in [MIL25a].
@@
Let $G$ be a finite $p$-group. The first nontrivial example of an endotrivial, and those which make the magic happen in the elementary abelian case, are as follows (c.f. [BG25, Definition 12.3]). If $p=2$, one has an endotrivial for $C_{2}$ given by $kC_{2}\to k,$ with $k$ in homological degree 0 and the nonzero differential given by the augmentation homomorphism. If $p$ is odd, one has an endotrivial for $C_{p}$ by truncating a periodic resolution of the trivial $kC_{p}$-module $k$, $kC_{p}\to kC_{p}\to k$ with $k$ in homological degree 0. For any subgroup $N\triangleleft G$ of index $p$, inflation yields the following complex
@@
Here $p$ is assumed odd, if $p=2$ then the degree 2 term is deleted. In the notation of Balmer–Gallauer, we set
$2^{\prime}:=\begin{cases}1&\text{if }p=2;\\ 2&\text{if }p>2;\end{cases}$
hence $u_{N}$ has length $2^{\prime}$. We keep the $u_{N}$ notation throughout the paper.
A more secluded example for $p=2$ and $G=D_{2^{n}}$ (here, we mean the dihedral group with $2^{n}$ elements, rather than the dihedral group acting on the regular polygon with $2^{n}$ edges) is as follows. Let $H_{1},H_{2}$ be nonconjugate, non-central subgroups of order 2. This choice is unique up to conjugacy and reordering. Then the following complex is endotrivial:
$kD_{2^{n}}\to k[D_{2^{n}}/H_{1}]\oplus k[D_{2^{n}}/H_{2}]\to k,$
with all maps between indecomposable permutation modules being augmentation homomorphisms, $k$ in homological degree 0.
Finally, we note a structural fact about endotrivial complexes; the trivial $kG$-module $k$ only appears up to homotopy once.
Proposition 2.8.
Let $G$ be a $p$-group and $C\in\operatorname{Pic}(\mathscr{K}(G))$ be an indecomposable endotrivial complex. Then $C_{h_{C}(G)}$ contains the trivial $kG$-module $k$ as a direct summand exactly once, and for all $j\neq h_{C}(G),$ $C_{j}$ does not contain $k$ as a direct summand.
Proof.
Since $\Psi^{G}(C)\cong k[h_{C}(G)]$, $C_{h_{C}(G)}$ must contain a direct summand isomorphic to $k$, and [MIL24, Lemma 5.8] with $\mathcal{Y}:=s_{p}(G)\setminus\operatorname{Syl}_{p}(G)$ proves the rest. ∎
2.1. Borel-Smith functions and representation spheres
Borel-Smith functions classically arise not in the context of representation theory, but in equivariant topology. To close out our preliminaries, we consider their origin: representation spheres, following [TOM87].
@@ 2.1. Borel-Smith functions and representation spheres
We first recall some important facts about Borel-Smith functions. Let $G$ be a group and $F$ a field. Given a $FG$-module $V$, the dimension function associated to $V$ is the superclass function
@@
If $F$ has characteristic 0, this induces a group homomorphism $R_{F}(G)\to\operatorname{CF}(G)$. When $F=\mathbb{R}$, it is common to write $\operatorname{RO}(G)=R_{\mathbb{R}}(G)$, where $\operatorname{RO}(G)$ denotes the Grothendieck ring of representation spheres, i.e., 1-point compactifications of real representations (see [TOM87, Page 12]). In this case, the dimension function is equivalently computing the dimension of the representation sphere $S^{V^{H}}=(S^{V})^{H}$ of $V^{H}$.
@@
[TOM87, Theorem 5.4 and Theorem 5.13, pages 211 and 216] Let $G$ be a nilpotent group. The image of $\dim\colon RO(G)\to\operatorname{CF}(G)$ is the group of Borel-Smith functions $\operatorname{CF}_{b}(G)$. Moreover, if $f$ is a effective Borel-Smith function returning non-negative values, there exists a real representation $V$ for which $\dim(V)=f$.
The following is immediate.
@@
Therefore, it suffices to work with effective endotrivials, which we will see are particularly well-behaved. In fact, when $G$ is a nilpotent group, the dimension function gives a canonical effective basis of $\operatorname{CF}_{b}(G)$, which arises from the real irreducible representations of $G$. Therefore, if $G$ is a finite $p$-group, we have a corresponding canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K})$ consisting of effective endotrivials.
@@
Since the image of the dimension homomorphism is precisely $\operatorname{CF}_{b}(G)$, it suffices to show that the set $\mathcal{B}:=\{f_{1},\dots,f_{n}\}$ is linearly independent after removing duplicates. [TOM87, Proposition 5.9, page 213] asserts that $\ker(\dim)$ is generated by elements of the form $V-\psi^{k}(V)$, where $V$ is an irreducible real representation, $\psi^{k}$ is the $k$-th Adams operation, and $k$ is coprime to $|G|$. Therefore, the duplicates in $\mathcal{B}$ arise from Adams-conjugates, hence any set of real irreducible representations with no Adams-conjugates will correspond to a linearly independent set of Borel-Smith functions. ∎
@@
We call the Borel-Smith functions (resp. endotrivials) associated to the real irreducible representations of $G$ the irreducible Borel-Smith functions (resp. endotrivials). These objects are necessarily effective.
On the other hand, an indecomposable endotrivial means what one would expect. It is an easy exercise to verify that in $\mathscr{K}(G)$, for every element $C\in\operatorname{Pic}(\mathscr{K}(G))$, there is a unique indecomposable representative of the isomorphism class of $C$.
@@
$\dim_{\mathbb{R}}V^{H}=\frac{1}{|H|}\sum_{h\in H}\chi(h)$
providing a computational method of computing the basis of $\operatorname{CF}_{b}(G)$.
Checking character tables of $G$ a finite $p$-group of normal $p$-rank one shows the basis of $\operatorname{Pic}(\mathscr{K}(G))$ obtained in [MIL24, Section 6] coincides with the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$. We remark that this is rather surprising, as the computations of [MIL24] were performed entirely ad-hoc, as they occurred well before the classification of $\operatorname{Pic}(\mathscr{K}(G))$.
Notation 2.15.
Let $G$ be a finite $p$-group. Given a $\mathbb{R}G$-module $V$, let $f_{V}$ and $C_{V}$ denote the corresponding effective Borel-Smith function and effective indecomposable endotrivial respectively. That is, $f_{V}:=\dim V$ and $C_{V}$ is the unique indecomposable endotrivial for which $h_{C_{V}}=\dim V$. Definition 3.6 will explain how to obtain $C_{V}$ from $S^{V}$ directly.